# FinTruthQA: A Benchmark Dataset for Evaluating the Quality of Financial Information Disclosure

### v1

@@ 4.2 Settings and metrics
Implementation details. For ML-based models, we utilized the standard scikit-learn [^6] https://scikit-learn.org/stable/ library. For all PLM-based methods, we uniformly used PyTorch and HuggingFace’s framework [^7] https://github.com/HuggingFace for downstream fine-tuning and evaluation. The hyperparameter settings of our experiments using PLM-based model can be found in Supplement Section E. The tasks of question identification and question relevance encode only the questions as input, while answer readability and answer relevance required encoding both the question and the answer. The sequence starts with a special classification token [CLS], and the question and answer are separated by a special token [SEP]. During the continued pre-training, we added two special tokens, [Q] and [A], into the vocabulary and prefixed them to the question and answer, respectively. This helps the PLMs better identify the components of each sentence. The continued pre-training was performed based on the UER framework (Zhao et al., 2019). All tests of GPT-4 were conducted by calling OpenAI’s API with version GPT-4-0613.
@@
For the first two binary classification tasks, we used accuracy, precision, recall, and F1-score metrics. For the last two multi-class classification tasks, we calculated both micro and macro F1-scores, along with an additional metric: Quadratic Weighted Kappa (QWK) (Dong et al., 2017).
@@ 4.3 Results and analysis
Tables  2,  3, and  4 present the performance of different methods in FinTruthQA. Experimental results reported with error bars can be found in Supplement Section F.
@@ 4.3.1 Task 1: Question identification
Question identification Question relevance
Method Model Accuracy F1 Precision Recall Accuracy F1 Precision Recall
ML-based TF-IDF + LR 0.9410 0.9688 0.9452 0.9935 0.9924 0.9962 0.9924 1.0000
TF-IDF + RF 0.9370 0.9667 0.9423 0.9924 0.9924 0.9962 0.9924 1.0000
TF-IDF + SVM 0.9390 0.9678 0.9424 0.9946 0.9924 0.9962 0.9924 1.0000
TF-IDF + KNN 0.9330 0.9645 0.9411 0.9891 0.9924 0.9962 0.9924 1.0000
PLM-based BERT 0.9617 0.9795 0.9735 0.9857 0.9983 0.9992 0.9983 1.0000
@@
Method Model F1 (Micro) F1 (Macro) F1 (Weighted) QWK
ML-based TF-IDF + LR 0.8135 0.3510 0.7878 0.4152
TF-IDF + RandomForest 0.7950 0.3156 0.7540 0.2912
TF-IDF + SVM 0.8113 0.3362 0.7763 0.3669
TF-IDF + KNN 0.7852 0.3015 0.7395 0.2418
PLM-based BERT 0.8550 0.4203 0.8429 0.6293
RoBERTa 0.8539 0.4035 0.8405 0.6267
FinBERT 0.8790 0.4189 0.8629 0.6889
Mengzi-fin 0.8648 0.4109 0.8495 0.6512
DKPLM 0.8713 0.4355 0.8572 0.6666
SBERT-nli 0.8451 0.3914 0.8274 0.5776
RoBERTa-extractive-qa 0.8648 0.4109 0.8495 0.6512
BERT (Large) 0.8582 0.4035 0.8407 0.6209
RoBERTa (Large) 0.8571 0.4064 0.8441 0.6387
Table 2 shows the results for task question identification. In this task, we evaluated the performance of various models of ML-based and PLM-based. Among the ML-based methods, the LR classifier showed the best performance with an accuracy of 94.10% and an F1 score of 96.88%. Meanwhile, the BERT outperformed all ML-based models, achieving an accuracy of 96.17%, an F1 score of 97.95%, a precision of 97.35%, and a recall of 98.57%. The results suggest that it not only correctly identifies a higher proportion of real questions (high recall) but also maintains a high level of correctness in its positive predictions (high precision). This balanced performance in precision and recall underscores the model’s capability to handle the nuances in question identification effectively. This may be attributed to the PLM’s superior ability to encode text compared to TF-IDF.
@@ 4.3.2 Task 2: Question relevance
The evaluation results for the task question relevance are detailed in Table 2. The PLM-based model outperformed the ML-based models across all metrics, achieving an accuracy of 99.83%, an F1 score of 99.92%, a precision of 99.83%, and a recall of 100%. This indicates that the model effectively identifies relevant questions while also correctly classifying the few negative samples, showcasing a balanced performance. In contrast, the TF-IDF-based models all demonstrated identical performance metrics and tended to predict all samples as positive. This is because the label distribution for this task is highly skewed (shown in Table 1), with 910 positive samples and only 7 negative samples in the test set. This imbalance poses a significant challenge for accurate classification, as models can easily achieve high recall by predicting all samples as positive, but this approach would fail to identify the few negative samples correctly.
@@ 4.3.3 Task 3: Answer readability
The evaluation results for the task answer readability are shown in Table 3. Among the ML-based models, the LR classifier exhibited the highest performance, achieving an F1 Micro score of 81.35% and a QWK of 0.4152. While LR provided reasonable performance, the relatively low F1 Macro score highlights difficulties in handling class imbalance effectively. In contrast, the PLM-based models demonstrated a substantial improvement over the ML-based models. The BERT achieved an F1 Micro score of 85.50% and a QWK of 0.6293, showcasing its ability to manage the complexity and variability in readability assessment. FinBERT outperformed all other models with an F1 Micro score of 87.90%, an F1 Macro score of 41.89%, an F1 Weighted score of 86.29%, and the highest QWK of 0.6889. FinBERT’s pre-training on financial text gave it a distinct advantage in this domain-specific task, allowing it to achieve the highest accuracy and alignment with human judgments. We also visualize the confusion matrix for our best-performing model (FinBERT) in this task in Figure 4(a). While the model performs well in identifying answers with the best readability, it needs significant improvement in distinguishing lower readability levels.
@@ 4.3.4 Task 4: Answer relevance
Method Model F1 (Micro) F1 (Macro) F1 (Weighted) QWK
ML-based TF-IDF + LR 0.7001 0.5032 0.6549 0.4117
TF-IDF + RF 0.6979 0.4366 0.6241 0.3627
TF-IDF + SVM 0.7056 0.4496 0.6339 0.3906
TF-IDF + KNN 0.6521 0.3863 0.5787 0.2322
PLM-based BERT 0.7797 0.6638 0.7638 0.6479
RoBERTa 0.7710 0.6503 0.7534 0.6177
FinBERT 0.7688 0.6710 0.7623 0.6496
Mengzi-fin 0.7710 0.6705 0.7635 0.6484
DKPLM 0.7612 0.6540 0.7541 0.6375
SBERT-nli 0.7426 0.5790 0.7086 0.5258
RoBERTa-extractive-qa 0.7841 0.6795 0.7732 0.6641
BERT (Large) 0.7721 0.6800 0.7664 0.6331
RoBERTa (Large) 0.7841 0.6916 0.7772 0.6565
PLM with CPT-based FinBERT 0.8026 0.6958 0.7886 0.6951
RoBERTa-extractive-qa 0.7786 0.6626 0.7662 0.6619
BERT (Large) 0.7722 0.6756 0.7639 0.6212
RoBERTa (Large) 0.7328 0.6195 0.7214 0.5372
LLM-based Direct score (GPT-4) 0.6739 0.6311 0.7014 0.5906
Score with CoT (GPT-4) 0.5976 0.5845 0.6428 0.5751
@@
Table 4 shows the experiment results on task answer relevance. TF-IDF + SVM performs the best among ML-based models, with an F1 Micro score of 70.56% and a QWK of 0.3906. Regarding PLM-based models, RoBERTa-extractive-qa stood out with an F1 Micro score of 78.41% and a QWK of 0.6641, demonstrating significantly better performance than ML-based models. This can be attributed to its pretraining on QA datasets, which likely enhanced its ability to understand and evaluate answer relevance effectively. According to the performance of the PLMs with continued pretraining presented in Table 4, further pertaining brings the most improvement on FinBERT, which was initially trained on financial corpora, achieving a Micro F1 score of 80.26% and 0.6951 QWK score. On the other hand, continued pretrained does not improve model performance on several other PLMs (e.g. RoBERTa-extractive-qa, BERT large, RoBERTa large). The results could be attributed to the fact that FinBERT’s initial training presumably provided it with a foundational understanding of the financial lexicon and contexts, which was further enriched through pre-training. Other models might suffer from a misalignment between the model’s learned representations and the domain-specific requirements. The results of GPT-4 with and without CoT are presented at the bottom of Table 4, GPT-4 models are largely behind PLM-based models, highlighting the challenge of state-of-the-art LLM on Chinese financial information disclosure understanding. Direct GPT-4 scoring yielded better results compared to GPT-4 with CoT. A potential reason is that GPT-4 was primarily trained on English corpora; it lacks sufficient financial domain knowledge in Chinese. This may result in knowledge gaps and hallucinations during the CoT process, thereby reducing the model’s performance. Another study also observed a similar phenomenon where CoT failed in healthcare domain prediction with Chinese text. (Liu et al., 2024) We anticipate that further domain adaptation methods, such as few-shot learning or retrieval-augmented generation (Lewis et al., 2020), may improve the performance of LLMs on this task. Figure 4(b) shows the confusion matrix for the best-performing model (FinBERT with continued pretraining). It demonstrates that the model effectively captures answers that are completely related to the questions (labeled as 3). However, the model struggles to distinguish partially unrelated answers (labeled as 2) from completely related ones, where the nuanced differences are subtle and challenging.
@@
Figure 4: Confusion matrix for the best performing model in Task 3 Answer readability and Task 4 Answer relevance.
@@ 5 Limitations
There are several limitations in this study. First, the Chinese cultural and regulatory context of SSE and SZSE may limit the generalizability of the findings to other markets. Future research is needed to explore the applicability of the developed models and methodologies in different contexts and validate the findings across diverse settings. Second, the annotation process inherently involves a degree of subjectivity, as differences in annotators’ interpretations can introduce variability in the labeled data. To mitigate this, we conducted multiple rounds of pre-annotation and formal annotation with professional annotators, along with thorough error analysis and expert reviews to ensure consistency. Third, while precision, recall, F1-score, and quadratic weighted kappa (QWK) are standard metrics for evaluating model performance, they may not fully capture the qualitative aspects of information disclosure quality. The reliance on these metrics could lead to an overemphasis on quantitative evaluation while neglecting interpretability. Lastly, the task of question relevance is highly imbalanced, and models on this task are not well compared. We decide to keep this task in our benchmark to reflect the real-world data distribution.
@@ 6 Conclusion
In this study, we introduced FinTruthQA, a benchmark designed to evaluate advanced NLP techniques for the automatic quality assessment of information disclosure in financial Q&A data. We created a dataset comprising 6,000 real-world financial Q&A entries, each manually annotated based on four conceptual dimensions of accounting. Experiments revealed the effectiveness of pre-trained BERT models in tackling these tasks. In particular, we observed that models pre-trained on financial corpora, as well as those designed for related tasks such as QA, outperformed other variants of BERT and GPT-4. This emphasized the importance of domain-specific and task-oriented training in improving performance. Although existing NLP models showed reasonable performance on this benchmark, there is still significant room for improvement, especially in the task of answer relevance, which calls for better NLP models. This benchmark can serve as a new text understanding for evaluating LLMs and it provides a novel, valuable, and practical sandbox for improving financial information disclosure with AI and data-driven approaches.
Checklist

### v2

@@ 4.2 Settings and metrics
Implementation details. For the ML-based models, we utilized the standard scikit-learn [^7] https://scikit-learn.org/stable/ library. For all PLM-based methods, we used PyTorch [^8] https://github.com/pytorch/pytorch and HuggingFace’s framework [^9] https://github.com/HuggingFace for downstream fine-tuning and evaluation. Detailed hyperparameter settings for our PLM-based experiments can be found in Appendix Section E. The tasks of question identification and question relevance only required encoding the questions as input, while answer readability and answer relevance required encoding both the question and the answer. The sequence started with a special classification token [CLS], and the question and answer were separated by the [SEP] token . During continued pre-training, we added two special tokens, [Q] and [A], into the vocabulary and prefixed them to the question and answer, respectively. This helps the PLMs better identify the components of each sentence. The continued pre-training was performed based on the UER framework (Zhao et al., 2019). All tests of LLMs were conducted using the transformer framework [^10] https://github.com/huggingface/transformers, except for the GPT-4 experiments, which were conducted by calling OpenAI’s API with version GPT-4-0613. Specifically, we used Llama-3.1 with 8 billion parameters, and both Qwen2 and Mistral with 7 billion parameters.
Category Model Question identification (%) Question relevance(%)
Accuracy Precision Recall F1 Accuracy Precision Recall F1
[…]
@@
For Question Identification and Question Relevance, which are binary classification tasks, we used accuracy, precision, recall, and F1-score as evaluation metrics. For Answer Readability and Answer Relevance, which are multi-class classification tasks, we calculated both micro and macro F1-scores, along with the Quadratic Weighted Kappa (QWK) (Dong et al., 2017), which is well-suited for ordinal classification by penalizing misclassifications based on their distance from the true label. Each model was independently run five times, and the mean and variance of the results are reported.
@@ 4.3.1 Task 1: Question identification
Category Model F1 (Micro) F1 (Macro) F1 (Weighted) QWK
[…]
In this task, we benchmarked several ML-based and PLM-based models, with their performances summarized in Table 2. The ML-based models showed consistent results with minor variations. Both TF-IDF + LR and TF-IDF + RF reached an accuracy of 93.70% and an F1 score of 96.66, with closely aligned precision and recall. TF-IDF + SVM exhibited slightly lower accuracy but achieved the highest recall among the ML models, demonstrating its strength in identifying relevant instances. TF-IDF + KNN, while performing well with an accuracy of 92.67%, had the lowest precision and F1 score in this group. On the other hand, the PLM-based BERT model outperformed all ML-based methods, achieving the highest accuracy, precision, and F1 score. This result highlights the PLM model’s ability to capture deeper contextual nuances, which traditional frequency-based methods like TF-IDF struggle to achieve.
@@ 4.3.2 Task 2: Question relevance
Table 2 presents the evaluation results for the question relevance task. The PLM-based model outperformed the ML-based models on all metrics except recall, achieving an accuracy of 99.67%, an F1 score of 99.85%, precision of 99.72%, and recall of 99.96%. This demonstrates its ability to effectively identify relevant questions while maintaining a strong balance in correctly classifying rare negative samples. In contrast, the TF-IDF-based models showed nearly identical performance across metrics, primarily predicting all samples as positive. Given the highly imbalanced label distribution (910 positive vs. 7 negative samples in the test set), these models achieve high recall by predicting positives but struggle to correctly identify the few negative instances, reflecting the challenge posed by the skewed data.
@@ 4.3.3 Task 3: Answer readability
The evaluation results for the task answer readability are shown in Table 3. Among the ML-based models, the LR classifier exhibited the highest performance, achieving an F1 Micro score of 81.72% and a QWK of 41.20%. While LR provided reasonable performance, the relatively low F1 Macro score highlights its difficulties in handling class imbalance effectively. In contrast, the PLM-based models demonstrated significant improvements over the ML-based models. For instance, BERT achieved an F1 Micro score of 86.04% and a QWK of 63.52%, showcasing its superior ability to capture the complexities of readability assessment through richer contextual understanding.
Detailed Evaluation of FinBERT’s Performance.
FinBERT outperformed all other models with an F1 Micro score of 87.57%, an F1 Macro score of 41.34%, an F1 Weighted score of 85.98%, and the highest QWK of 66.53%. Its pre-training on financial text provided a distinct advantage for this domain-specific task, enabling it to achieve the highest accuracy and alignment with human judgments. To further analyze FinBERT’s performance, we visualize its confusion matrix in Figure 4(a). We find that the model performs well in recognizing answers with the highest readability, but it struggles to accurately differentiate between the lower readability levels, highlighting the need for improvement in capturing finer distinctions in readability.
@@ 4.3.4 Task 4: Answer relevance
Category Model F1 (Micro) F1 (Macro) F1 (Weighted) QWK
[…]
LLM-based GPT-4 (Direct score) 67.39 63.11 70.14 59.06
GPT-4 (with CoT) 59.76 58.45 64.28 57.51
Llama-3.1 (Direct score) 53.00 41.59 56.13 43.24
Llama-3.1 (with CoT) 46.46 40.82 51.57 39.56
Qwen2 (Direct score) 60.52 31.33 52.83 23.68
Qwen2 (with CoT) 59.00 38.97 57.36 42.35
Mistral (Direct score) 54.53 30.40 50.99 29.57
Mistral (with CoT) 59.32 23.03 60.20 38.90
PLM-based Methods Consistently Outperform ML-based ones.
As shown in Table 4, all PLM-based methods can achieve better performance than ML-based methods across all metrics. For instance, RoBERTa-extractive-qa achieved the highest Micro F1 score of 78.33%, while Mengzi-fin excelled with the highest Macro F1 score of 68.96% and a strong QWK of 66.31%. Their superior performance can be attributed to pre-training on task-specific and domain-specific data: RoBERTa-extractive-qa was pre-trained on QA datasets, which enhanced its ability to comprehend and evaluate answer relevance, while Mengzi-fin was pre-trained on financial corpora, enabling it to better capture domain-specific nuances in the financial context.
Impact of Continued Pre-training on PLM-based Methods.
We selected four PLM-based models and further pre-trained them using 900,000 Q&A pairs. As shown in Table 4, we observed that only FinBERT demonstrated significant performance improvements, achieving the highest Micro F1 score of 80.04% and a QWK of 67.89%, indicating its enhanced ability to assess answer relevance in financial contexts. In contrast, the other three models (RoBERTa-extractive-qa, BERT Large, and RoBERTa Large) did not exhibit consistent improvements. This suggests that domain alignment plays a crucial role in the effectiveness of continued pre-training, as FinBERT’s domain-specific pre-training on financial texts provided a strong foundation, whereas the other models may have experienced a mismatch between their learned representations and the specific requirements of the task, resulting in less notable gains.
Suboptimal Performance of LLMs in Chinese Financial Contexts.
@@
As shown in Table 4, we also explored the performance of some representative LLMs on answer relevance task. We observe that GPT-4 (Direct score) consistently outperform other LLMs but still lagged behind supervised PLM-based models, underscoring the difficulty LLMs face in understanding Chinese financial disclosures. Notably, GPT-4 with CoT did not outperform GPT-4 Direct score, suggesting that the CoT mechanism may introduce confusion rather than improving performance in this task. A potential reason is that GPT-4 was primarily trained on English corpora; it lacks sufficient financial domain knowledge in Chinese. This may result in knowledge gaps and hallucinations during the CoT process, thereby reducing the model’s performance. Another study also observed a similar phenomenon where CoT failed in healthcare domain prediction with Chinese text. (Liu et al., 2024) We also tested Chinese financial LLMs, including CFGPT2(Li et al., 2023) and DISC-FinLLM(Chen et al., 2023), but found that these models could hardly follow instructions and failed to generate reasonable predictions. As a result, we decided to exclude them from our study. We anticipate that further domain adaptation methods, such as few-shot learning or retrieval-augmented generation (Lewis et al., 2020), may improve the performance of LLMs on this task. Figure 4(b) shows the confusion matrix for the best-performing model (FinBERT with continued pretraining). It demonstrates that the model effectively captures answers that are completely related to the questions (labeled as 3). However, the model struggles to distinguish partially unrelated answers (labeled as 2) from completely related ones, where the nuanced differences are subtle and challenging. The confusion matrices for GPT-4 (direct score) are shown in Figures 4(c), which illustrate that the model tends to mix up answers that are completely irrelevant (labeled as 1) with those that are partially relevant (labeled as 2). Also, there’s a tendency for the model to predict label 2.
Error Analysis on GPT-4 Performance.
We have also conducted a thorough examination of bad cases to better understand the prevalent error types exhibited by GPT-4 in relevance scoring. One common error we identified occurs when the model generates an answer that revolves around the key terms of the query but fails to address the core question, leading to an incorrect judgment of relevance where there is none. Additionally, issues with pronoun resolution have led to errors where the question concerns one subject, such as employee benefits, but the answer discusses another, such as salaries, resulting in a misjudgment of irrelevance. Another notable issue is the model’s handling of answers that state ”the situation is unclear.” In such cases, GPT-4 often misclassifies these as unclear responses, rather than recognizing them as valid.
@@
(c) GPT-4 (direct score) for answer relevance task.
Figure 4: Confusion matrix for models in Task 3 Answer readability and Task 4 Answer relevance.
{CJK}
UTF8gbsn
Table 5: Examples of GPT-4 (direct score) bad cases and error analysis.
Question Answer Prediction Label Error analysis
请问贵公司2020年丰水季节弃电量是多少? 尊敬的投资者您好!公司丰水期弃水电量受当年气候、来水量、网络负荷、相关政策等综合影响。谢谢关注。 2 1 Model focuses on matching keywords instead of understanding the core intent of the question.
Could you please tell me what the amount of curtailed electricity was for your company during the 2020 flood season? Dear investors, hello! The amount of abandoned hydroelectric power during the company’s flood season is influenced by various factors, including climate, water inflow, network load, and relevant policies. Thank you for your attention.
董秘你好,贵公司今年利润下降,有没有考虑过降低员工待遇或者取消年终奖来维持利润 您好!公司注重提升员工的幸福感,在2017年初,公司普遍上调员工工资。对于年终奖,公司将综合考虑员工考核结果及公司业绩等多方面因素来确定。 2 3 Error in pronoun resolution (e.g., salaries vs. employee benefits), causing confusion between related subjects.
Hello, Company Secretary. Given your company’s decline in profits this year, have you considered reducing employee benefits or canceling year-end bonuses to maintain profits? Hello! The company places great emphasis on enhancing employee well-being. At the beginning of 2017, the company generally increased employee salaries. For the year-end bonus, the company will consider various factors, including employee performance evaluations and overall company performance, to determine the amount.
金风科技投资温州深远海海上风电项目拟引进塔筒,机舱罩等配套企业,公司是否参与 尊敬的投资者,该情况尚不明确。感谢您的建议。 2 3 Model fails to interpret ”unclear” as a valid state rather than an irrelevant response.
The investment by Jinfeng Technology in the Wenzhou Deep Sea Offshore Wind Power Project plans to introduce supporting enterprises for tower tubes, nacelle covers, etc. Is the company involved? Dear Investor, the situation is still unclear. Thank you for your suggestion.
@@ 5 Limitations
There are several limitations in this study. First, the Chinese cultural and regulatory context of SSE and SZSE may limit the generalizability of the findings to other markets. Future research is needed to explore the applicability of the developed models and methodologies in different contexts and validate the findings across diverse settings. Second, the annotation process inherently involves a degree of subjectivity, as differences in annotators’ interpretations can introduce variability in the labeled data. To mitigate this, we conducted multiple rounds of pre-annotation and formal annotation with professional annotators, along with thorough error analysis and expert reviews to ensure consistency. Third, while precision, recall, F1-score, and quadratic weighted kappa (QWK) are standard metrics for evaluating model performance, they may not fully capture the qualitative aspects of information disclosure quality. The reliance on these metrics could lead to an overemphasis on quantitative evaluation while neglecting interpretability. Lastly, the question relevance task suffers from a highly imbalanced dataset, which may limit the model’s ability to effectively learn from and evaluate rare instances. However, we chose to retain this task in the benchmark to reflect the real-world data distribution in financial information disclosure.
@@ 6 Conclusion
In this study, we introduce FinTruthQA, a benchmark designed to evaluate advanced NLP techniques for the automatic quality assessment of information disclosure in financial Q&A data. We constructed a dataset of 6,000 real-world financial Q&A entries, each manually annotated based on four key evaluation criteria of accounting. Experiments revealed the effectiveness of pre-trained BERT models in tackling these tasks. In particular, we observed that models pre-trained on financial corpora, as well as those designed for related tasks such as QA, outperformed other variants of BERT and GPT-4. This emphasized the importance of domain-specific and task-oriented training in improving performance. Although existing NLP models showed reasonable performance on this benchmark, there is still significant room for improvement, especially in the task of answer relevance, which calls for better NLP models. Our benchmark is not only useful for post-hoc regulatory assessments but also holds significant value in real-time applications. For instance, auditors, financial analysts, and corporate communication teams can leverage our framework for ongoing monitoring and assessment of information disclosure quality. This real-time application enhances market efficiency by enabling quicker identification and correction of inadequate disclosures, thereby benefiting all market participants, including retail investors. Moreover, our benchmark can serve as a valuable resource for academic research in natural language processing, particularly in areas concerning the automated assessment of text quality.
Appendix A Ethics
General ethical conduct.
Our research on assessing the quality of information disclosure in financial Q&A using BERT-based models utilizes publicly available data from the Shanghai Stock Exchange and the Shenzhen Stock Exchange’s interactive platforms. Following ethical guidelines, we ensured all data was anonymized to protect individual identities. The dataset consists of Q&A entries that are openly displayed on the interactive platforms and are used in compliance with the respective exchanges’ policies and the China Securities Regulatory Commission’s guidelines. To ensure transparency and reproducibility, we provide the code and pre-trained models used in our experiments, subject to any legal or ethical restrictions. To mitigate algorithmic bias, our annotation process included multiple rounds of pre-annotation and formal annotation involving professional annotators, ensuring a diverse and representative sample of Q&A data. Rigorous data cleaning and error analysis were conducted to address biases. We encourage users to perform their own checks to avoid potential biases related to specific sub-populations or other contextual factors.
Potential negative societal impacts.
The primary goal is to develop models for evaluating disclosure quality rather than conducting direct causal analyses. Users should be cautious when applying these models in contexts other than financial Q&A, as the models may have learned information specific to the financial domain and Chinese market settings. Additionally, over-reliance on automated systems can lead to the neglect of human judgment and expertise. We recommend using our models as supplementary tools to assist human analysts rather than replace them entirely. Although our work aims to improve transparency and information quality in financial markets, users should take appropriate precautions to avoid any potential negative societal impacts resulting from misapplication of the models.
Appendix B Annotation process
Figure 5: Guideline of the annotation process.
The guideline of the general annotation process is depicted in Figure 5. The annotation process consists of the following steps: